OpenStudy (anonymous):

Find dy/dx using implicit differentiation: xe^y=y-1

OpenStudy (amistre64):

..imply it then :) let y be some function of x, and apply the chain rule to the usual derivative rules

OpenStudy (amistre64):

i see the basics: product rule, e rule, power rule and constant rule

OpenStudy (anonymous):

so d/dx (xe^y) = d/dx (y-1) to start it off?

OpenStudy (amistre64):

yep, thats a fine start if you like that kind of notation

OpenStudy (anonymous):

I know d/dx = y-1 = ydx/dy im kind of stuck on d/dx(xe^y)

OpenStudy (amistre64):

d/dx (y - 1) d/dx (y) - d/dx (1) dy/dx - 0

OpenStudy (amistre64):

d/dx (xe^y) ; is a product rule at heart d/dx (x) e^y + x d/dx (e^y) dx/dx e^y + x dy/dx e^y e^y + x dy/dx e^y

OpenStudy (amistre64):

if you view this as the d/dx "absorbing" the variable; it gets easier to correlate

OpenStudy (amistre64):

\[\frac{d}{dx}(e^y)\] absorb the y, and run a normal e^x outcome \[\frac{dy}{dx}~e^y\]

OpenStudy (anonymous):

Oh so i got e^y/1-xe^y

OpenStudy (amistre64):

maybe, let me try it my way: \[x~ e^y = y - 1\] \[e^y+x~ y'e^y = y'\] \[e^y = y'-x~y'e^y\] \[e^y = y'(1-x~e^y)\] \[y'=\frac{e^y}{1-x~e^y}\]

OpenStudy (ranga):

You can also think of d/dx(e^y) as d/dy(e^y)dy/dx = (e^y)(dy/dx)

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

thanks a lot

OpenStudy (amistre64):

good luck

OpenStudy (amistre64):

ranga, i cant read thru that notation to well to comment :)

OpenStudy (amistre64):

i think i see it now ...

OpenStudy (ranga):

\[\frac{ d }{ dx}e ^{y} = \frac{ d }{ dy }e ^{y}\frac{ dy }{ dx } = e ^{y}\frac{ dy }{ dx }\]